OpenStudy (anonymous):

A 12-foot ladder is leaning against a wall. The bottom of the ladder is 5 feet away from the bottom of the wall. Approximately how high up the wall does the top of the ladder reach? A. 2.4 feet B. 7.0 feet C. 13.0 feet D. 10.9 feet

OpenStudy (anonymous):



OpenStudy (er.mohd.amir):

it make right angle triangle so use Protagoras theorem 12*12=5*5+H*H H*H=144-25 =119 H=10.9

OpenStudy (anonymous):

thats what i was thinking but thank you so much